OpenStudy (anonymous):

.What is the sum of a 6-term geometric series if the first term is 22 and the last term is 1,299,078? 1,395,030 1,461,460 1,527,890 1,594,320

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well you need to common ratio... before you find the sum. the formula for a term is \[a_{n} = a_{1} \times r^{n - 1}\] so using what you know \[1299078 = 22 \times r^5\] you need to solve for r. when you get r. the sum of a geometric series is \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] subsitute your values, n = 6, a = 22 and your r value to get the sum

OpenStudy (anonymous):

Whoa, Whaat ;-;